OpenStudy (anonymous):

Find the interval of convergence of \[\sum_{1}^{\infty} \frac{ 8x^n }{ n }\] I am not sure what I am doing wrong here, I'm pretty sure this should be the interval [-1/8,1/8) but that seems to be wrong.

OpenStudy (anonymous):

I have |x|<1/8 when I solve this, but apparently that isnt correct, unless my brackets are off.

OpenStudy (anonymous):

you are getting confused by the \(8\) ignore it

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ 8x^n }{ n }=8\sum_{n=1}^{\infty} \frac{ x^n }{ n }\]

OpenStudy (anonymous):

it is a constant, it is not being raised to the power of \(n\) so just pull it out front, it does not effect the interval of convergence

OpenStudy (anonymous):

ah crap, thanks, thats a silly mistake.

OpenStudy (anonymous):

(-1,1) with open brackets then

OpenStudy (anonymous):

no, i think you are good at \(-1\)

OpenStudy (anonymous):

i.e. \[[-1,1)\]

OpenStudy (anonymous):

ah right I see that, -1 is fine as it would end as - (-1). THanks for the help!

OpenStudy (anonymous):

yw